Purple Company records $200,000 in net income for 2019 before deducting any compensation or other payment to its sole owner, Kirsten. Kirsten is single and she claims the $12,200 standard deduction. Purple Company is Kirsten's only source of income.

Ignoring any employment tax considerations, compute Kirsten's after-tax income for each of the following situations.

Click here to access the 2019 individual tax rate schedule to use for this problem. Assume the corporate tax rate is 21%.

When required, carryout intermediate tax computations to the nearest cent and then round your final tax liability to the nearest dollar.

a. If Purple is a proprietorship. Kirsten withdraws $50,000 from the business during the year; she claims a $37,560 deduction for qualified business income. Kirsten's taxable income is
$, and her after-tax income is $.

b. Purple is a C corporation and pays out all of its after-tax income as a dividend to Kirsten.

Note: Individual taxpayers received preferential treatment regarding the taxation of qualified dividends (0%,15%,20%). For single taxpayers, the 0 percent rate applies to the first $38,600 of taxable income.

Purple Corporation's after-tax income is $ and Kristen's after tax income is $.

c. Purple is a C corporation and pays Kirsten a salary of $158,000. Kirsten's after-tax income is
$.
Answer #1

Part A

Income from proprietorship

$200,000

Less: deduction for qualified business income

(37560)

Less: standard Deduction

(12200)

Taxable income

150240

Tax on $150240

30232 (14382.50 + (24%*(150240-84200)))

After-tax income ($200,000 – $30232)

$169767

Part B

Tax on corporation’s net income of $200,000 = 200000*21% = 42000

Corporation’s after-tax income = 200000-42000 = 158000

Kirsten’s taxable income = 158000-12200 standard deduction = 145800

Kirsten’s tax on $146000 at rates applicable to dividends = (39375*0%)+((145800-39375)*15%) = 15964

Kirsten’s after-tax income = 158000-15964 = $142036

Part C

Taxable income of corporation = 200000-158000 = 42000

Taxable income of Kristen = 158000-12200 standard deduction = 145800

Kristen’s tax = (14382.50 + (24%*(145800-84200)))= 29167

Kirsten’s after-tax income = 145800-29167 = $116633